OpenStudy (anonymous):

Find the equation of a line with slope = 6 and passing through A(4, 5).

OpenStudy (adunb8):

dont you use y = mx+b?

OpenStudy (anonymous):

yes but how do you find b?

OpenStudy (anonymous):

i know m=6

OpenStudy (adunb8):

b is the y- intercept

OpenStudy (anonymous):

i know..how do you get that from the given info though?

OpenStudy (anonymous):

a. y = 6x + 19 b. y = 6x - 19 c. y = 6x - 20 d. y = 6x - 21

OpenStudy (anonymous):

your point is (x,y)! use that for the other variables so now b is the only one left to solve for :D

OpenStudy (anonymous):

those are the choices

OpenStudy (anonymous):

what do you mean @nj1202 ?

OpenStudy (anonymous):